OpenStudy (anonymous):

64b^2=16 solve equation by finding square roots

OpenStudy (anonymous):

64b^2-16=0 16(4b^2-1)=0 4b^2-1=0 (2b-1)(2b+1)=0 2b-1 = 0 and 2b+1 = 0 2b = 1 and 2b = -1 b = 1/2 and -1/2